how do i slove this problem
h(x)= x^2+4x-5 ; h(x)>0

Answer by richwmiller(9143)

Oh good, I haven't "sloved" a problem in a long time.
h(x)= x^2+4x-5
h(x) = (x-1)(x+5),h(x)>0
-5>x>1
Test a value of x in three ranges.
x<-5 ok -6
(-6)^2+4*(-6)-5>0
36-24-5=7
1>x>-5 not >0 test 0 h(x)=-5
and x>1 ok
2^2+4*2-5=4+8-5=7